OpenStudy (anonymous):

How are the functions y = x and y = x – 3 related? How are their graphs related? Can someone explain this to me

OpenStudy (anonymous):

Each output for y = x – 3 is 3 more than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated down 3 units. Each output for y = x – 3 is 3 more than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated up 3 units. Each output for y = x – 3 is 3 less than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated down 3 units. Each output for y = x – 3 is 3 less than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated up 3 units.

OpenStudy (anonymous):

y = x and y = x – 3 are these related do to same slope

OpenStudy (anonymous):

They do have the same slope just different y's the first one passes at (0,0)and the second at (0,-3)

OpenStudy (zaarin):



OpenStudy (anonymous):

that would be -x+3 wouldn't it

OpenStudy (anonymous):

@RavenLynette what your saying is that c is the right answer Each output for y = x – 3 is 3 less than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated down 3 units. "y's the first one passes at (0,0)and the second at (0,-3)" Right?

OpenStudy (anonymous):

Right :)